Summary

The seeing is bad. Most of

The temperature of the telescope is unstable, making the focus change a lot and it is very hard to converge. The red channel can not work. we are in blue-only mode. The second half of the night is more productive than the first half. We got one field on IC 348 (but we do not know where the data are); Segur-0 and Segur-3 in both B and r bands and Bootes12, 52, 59, 16, 64, 46 in U_spec

Details

Observers: Zoltan Bolag, Fuyan Bian

Supporting Astronomer: Marco Pedani

Operator: Wayne Wack

All Times UT

Forth night of the AZ queue.

2320 The sky looks good, though there are a lot of clouds around us. The region around zenith is very clear. The snow on the roof is melting.

2330 We are taking some bias frames. There are some horizontal lines on the red channel bias frames. We will restart the LBC and re-take some bias. Meet the 'CCD electronics temperature above safety limit on red channel' problem

0000 We restart the LBC.

0104 We are opening

0115 We re-take the bias frames, the red channel bias is coming out very slowly. Meet the some problem. We have no idea no this

0130 We are in Blue only mode, and we are taking flat_r (15s),

0145 OB execution ERROR: Blue channel filter problem, we restart the LBC. The red channel has the same overheating

0220 Go to the first scientific field IC 348, seeing is very bad, mirror is warm. The focus is hard to converge.

0236 The source is too close to the zenith. We slew to Aries to check the focus.

0336 We finally figure out why the focus does not converge. The operator adjust the focus for the right (wrong!) mirror, We should change the focus on left mirror (blue side)...

0339 We move back to IC 348.

0410 Finally, we get the focus converge. We take the scientific data IC348 in r band, Meet some data transfer problems, we may miss all the data.

0425 We restart the bino-mode, and take some test image to see whether the red channel is OK. And red channel electronics overhearting problem may be caused by the readout process. So for scientific fields, the red channel should be OK.

0545 The red channel shuttle can not be open...:(. and we go to Segue_0. Focus seeing estimate 1.16"

0600 We are taking scientific data for Segue_0 for B band. The first image has a bad tracking, the guider choose a galaxy, and the same as the second one We have to stop

0624 Do focus again. And add 20' to the Dec for Segue_0 in B band and taking data again. The first image looks good. The seeing is ~1.2

0715 Do focus and take Segue_0 in r band. The seeing is ~1.0. And the seeing becomes bad, The seeing for the fourth image is ~2", the last two image is out of focus.

0749 Go to Segue_3 in B and do focus. The seeing is 1.6-1.8 for the first four image and the fifth one become really bad greater than 2.0", and we have to stop to adjust the focus again. The wind reached 15 m/s

0832 Go to Segue_3 in r. The seeing is 1,0-1.2. The focus becomes very bad for the fifth one. The telescope temperature changes a lot. The focus can not be stable beyond 20 min

0900 Go to Bootes46 in U_spec, The seeing is bad, and the focus is hard to converge. The seeing is very bad, we have to stop to redo the focus.

0942 We restart Bootes46 again still very bad.

1033 We finally figure out the reason that U_spec is out of focus. When the telescope apply the difference of the focus between V and U band. The Z meet its upper limit. We apply the Z offset to -6 to make the focus work. We are taking Bootes46. The seeing is 1.5-1.8".

1054 We move to Bootes59 in U_spec and do the focus. the seeing is 1.6-1.8"

11:25 We move to Bootes19 in U_spec and do the focus. The seeing is 1.5-1.7"

11:54 We move to Bootes52 in U_spec and do the focus. The seeing is 1.3-1.5"

11:13 We move to Bootes12 in U_spec and do the focus. The seeing is 1.3-1.4"

11:52 We move to Bootes64 in U_spec and do the focus. The seeing is 1.0-1.2". The background for the last one lbcb.20090125.131458.fits is a little larger around 3000 ADU.

13:22 We take 5 bias frames, then another 5 bias frames

13:36 take r, R, U flats.

A quick look of the image quality is attached, if the Ellip is greater than 0.1. The image is thought to be elongated.

-- XiaohuiFan - 24 Jan 2009
Topic revision: r3 - 26 Jan 2009, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback